OpenStudy (anonymous):

find the slope of the following line. 5y=x+5

OpenStudy (mathmale):

Hello! The usual way to determine the slope of this straight line is to isolate y. Please do this by dividing both sides of the equation by 5.

OpenStudy (mathmale):

What do you get?

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

sorry 1/5

OpenStudy (mathmale):

If 5y=x+5 and we divide both sides by 5, we get 5y + x = 5 ---- -- --- 5 5 5 WOULD YOU please write out the whole equation, not just the (1/5)?

OpenStudy (mathmale):

y = ??

OpenStudy (mathmale):

Ms. Menendez: still there?

OpenStudy (anonymous):

y=5/5

OpenStudy (anonymous):

y=1

OpenStudy (mathmale):

What I was trying to explain was that you should start with the given equation, 5y=x+5, and divide it through by 5 to obtain a formula for y. That formula is y=(1/5)x + 1. Then, look at the coefficient in front of the x. What is this coefficient?

OpenStudy (anonymous):

1/5

OpenStudy (mathmale):

Good. That's your slope, and that's the proper way to find it. I'm sorry this took so long, but I'd rather that you understand where that slope comes from and why.

OpenStudy (mathmale):

So, after you've solved 5y=x+5 for y, you merely take the coefficient of x as your slope, m. OK?

OpenStudy (anonymous):

okay i got it

OpenStudy (mathmale):

Good luck! See you. Bye for now.

OpenStudy (anonymous):

okay bye see you soon :)